There are three consecutive intergers. The sum of the first and the third is equal to twice the second
Answer by checkley79(3341)

Let x be the first integer
Let x+1 be the second integer
Let x+2 be the third integer.
x+(x+2)=2(x+1)
2x+2=2x+2
This tells me that ANY series of 3 consecutive integers will meet these requirements.
Examples:
4,5,6
4+6=2*5
10=10
55,56,57
55+57=2*56
112=112
451,452,453
451+453=2*452
904=904